OpenStudy (anonymous):

I think it is c... can I get a verification? Find the ordered pairs for the x- and y-intercepts of the equation 3x − 2y = 18 and select the appropriate option below. a.The x-intercept is (0, 6); the y-intercept is (−9, 0). b.The x-intercept is (6, 0); the y-intercept is (0, −9). c.The x-intercept is (3, 0); the y-intercept is (0, −2). d.The x-intercept is (0, 3); the y-intercept

OpenStudy (hyperlunatic):

hi

OpenStudy (hyperlunatic):

will help

OpenStudy (kenshin):

3x-18=2y, (3/2) x - 9 = y, so -9 is the y intercept, when y=0 (3/2)x = 9, x=6 so it's b

OpenStudy (anonymous):

Good job Kenshin.

OpenStudy (anonymous):

Thank you!